Solve by following systems by substitution
X+2y=1
X=y-5

Answer by mananth(16946)

X+2y=1
X=y-5
(y-5)+2y=1
3y-5=1
3y=6
y=2
plug y in (2)
x=y-5
=2-5
=-3
x=-3